Series connection solve for the total capacitance, total charge and individual voltage.


C1 = 3.5 x 10^-12

C2 = 2.4 x 10^-12

C3 = 1.13 x 10^-12

V = 1700 V


1
Expert's answer
2021-10-31T18:13:15-0400

"C=\frac{C_1C_2C_3}{C_1C_2+C_2C_3+C_2C_3}=0.63\cdot 10^{-12}~F,"

"q=CU=1.07\cdot 10^{-9}~C,"

"U_1=\frac q{C_1}=306~V,"

"U_2=\frac q{C_2}=446~V,"

"U_3=\frac q{C_3}=948~V."